OpenStudy (anonymous):

Which of the following is equal to the square root of the cube root of 5 ?

OpenStudy (anonymous):

5 to the power of 1 over 3 5 to the power of 1 over 6 5 to the power of 2 over 3 5 to the power of 3 over 2

OpenStudy (anonymous):

im not sure

jabez177 (jabez177):

Try and input it on a calculator. See what you get. :P

OpenStudy (anonymous):

so @Mehek14 ?

OpenStudy (anonymous):

so x is 5 and b is 3

OpenStudy (anonymous):

and a is 1

jabez177 (jabez177):

http://prntscr.com/99ab2i http://prntscr.com/99abc0 http://prntscr.com/99abjj http://prntscr.com/99abqx This is the question so now which one of these equal to this? http://prntscr.com/99ac0d

OpenStudy (anonymous):

thanks @Mehek14 and @jabez177

OpenStudy (anonymous):

1

jabez177 (jabez177):

Right on!

OpenStudy (anonymous):

thanks:D

OpenStudy (anonymous):

\[\sqrt{\sqrt[3]{5}}=\sqrt[6]{5} \]

OpenStudy (anonymous):

\[\sqrt[6]{5}=1.30766 \]